Pleas help me solve this equation: -3=-a/13
Answer by JulietG(1812)

-3=-a/13
Let's get rid of the fraction first.
Multiply each side by 13
-3*13 = -a/13 * 13
-39 = -a
Multiply each side by -1
-39 * -1 = -a * -1
39 = a